OpenStudy (anonymous):

A box contains 95 pink rubber bands and 90 brown rubber bands. You select a rubber band at random from the box. Find each probability. write the probability as a fraction in simplest form. a. find the theoretical probability of selecting a pink rubber band. b. find the theoretical probability of selecting a brown rubber band. P.S. it would be really helpful if you can explain how you got the answer or even work it out, Thank you to anyone who helps me out!

OpenStudy (anonymous):

P( a pink rubber) = 95/(95+90) = 95/185 = 19/37

OpenStudy (anonymous):

Thank you sooooooo much @samman8801!!!!

OpenStudy (anonymous):

Np, I hope you understood what I put, if you ever need me again just msg me :D

OpenStudy (anonymous):

Okay thanks again! @samman8801

OpenStudy (anonymous):

No problem